# https://sinhasaptarshi.github.io/escounts

### v1

@@ 3.3 Time-Shift Augmentations
The predicted density map $\tilde{\mathbf{d}}$ results from encoding a video with $\mathcal{E}$, over non-overlapping sliding windows. However, as each window is of fixed temporal resolution, repetitions may span over multiple windows. Thus, we include time-shifting augmentations in which the start time of the encoded video is adjusted to allow for different spatiotemporal tokens. We train with augmentations of the start time whilst at inference, we use an ensemble of time-shift augmentations for a more robust estimation. We use multiple overlapping sequences as shown in Figure 4 and combine the predicted density maps over $\mathbf{K}$ shifted start/end positions. We obtain the final predicted density map by temporally aligning and averaging the predictions.
$\tilde{\mathbf{d}}_{t}=\frac{1}{|\mathbf{K}|}\underset{k\in\mathbf{K}}{\sum}\tilde{\mathbf{d}}^{k}_{t+\epsilon_{k}}$ (7)
where $\epsilon_{k}$ is the shifting for each $k\in\mathbf{K}$.
@@ 4 Experiments
We overview the used datasets, implementation details, and evaluation metrics in Sec. 4.1. We include quantitative and qualitative comparisons to state-of-the-art methods in Sec. 4.2. We ablate over different ESCounts settings in Sec. 4.3. We demonstrate ESCounts’ ability to utilise exemplars at inference, in Sec. 4.4.
@@ 4.1 Experimental Setup
RepCount [20] contains videos of workout activities of varying repetition durations and appearances. Annotations include count labels alongside start and end timestamps per repetition. We use the publicly available Part-A of the dataset with $758$, $131$, and $152$ videos for train, val, and test respectively. We tune the hyperparameters on the val set and report our results on the test set.
@@
Countix [16] is a subset of Kinetics [9] containing videos of repetitive actions and their associated counts. It is the largest VRC dataset with $4588$, $1450$, and $2719$ videos for train, val, and test respectively. Count annotations are provided per video, without individual repetition start-end times. We use pseudo-repetition annotations by dividing the videos into uniform segments based on the ground truth count, as Countix does not have many pauses or interruptions between counts. We use the pseudo-labels for estimating the density maps, without additional annotations, to compare directly to other methods.
@@
UCFRep [65] is a subset of UCF-101 [51] consisting of 420 videos for train and 106 videos for val from 23 categories. Following [65, 30], we report our results on the val split as no test set is available. Like RepCount, it includes annotations of start and end timestamps for each repetition.
@@
Implementation Details. Unless specified otherwise, we use VideoMAE-VIT-B [55] as our encoder $\mathcal{E}$ with weights pre-trained on Kinetics-400 [9]. We sample frames from variable-length videos every 4 frames using a sliding window of $64$ frames. At each window, our encoder’s input is of $16\times 224\times 224$ size, and the output is $8\times 14\times 14$, resulting in $1568$ spatiotemporal tokens. We use $C=512$ channels333when encoders have a different output, we add a fully connected layer to map to $C$. The input to the decoder is of variable length $M=1568\frac{\mathcal{R}}{64}$, where $\mathcal{R}$ is the total number of frames in the video at raw framerate. Exemplars are sampled uniformly with $16$ frames between the start and end of a repetition.
@@
The encoder is frozen with only the decoder and the zero-shot latent $\mathbf{z}_{0}$ being trained. We use $L=2$ and $L^{\prime}=3$. We ablate $L$ and $L^{\prime}$ in the supplementary material. We train our model on a single Tesla V100 for 300 epochs with a batch size of 1, to deal with variable-length videos, and accumulate gradients over 8 batches. We use an initial learning rate of $55e-$ 5 and decay it every $60$ epochs by $0.8$, and a weight decay of $25e-$ 2. For each training instance, we randomly set the number of exemplars $|\mathcal{S}|\sim\{0,1,2\}$ and sample $\mathcal{S}$ exemplars. We set the chance of sampling exemplars from a different video to $p=0.4$.
@@
At inference, we only use the learnt latent to predict repetition counts. We aggregate predictions over $|\mathbf{K}|=4$ sequences.
@@
Evaluation Metrics. Following previous repetition counting works [20, 16, 67], we use Mean Absolute Error (MAE) and Off-By-One accuracy (OBO) as evaluation metrics, calculated as Eqs. 8 and 9 respectively. Inspired by image counting methods [10, 3], we introduce Root-Mean-Square-Error (RMSE) in Eq. 10 for VRC providing a more robust metric for diverse counts compared to MAE’s bias towards small counts. With continuous improvements in repetition counting methods, we also report the off-by-zero accuracy (OBZ) in Eq. 11 as a tighter metric than the corresponding OBO for precise counts.
@@ 4.2 Comparison with State-of-the-art
In Tab. 2, we compare our method, ESCounts, to prior methods on the three datasets. We also provide results on the same backbone as the best-performing method on each dataset, for fair and direct comparison to previous works.
Table 1: Comparison of VRC methods. $\dagger$ represents multi-modal models that use added audio or flow. denotes results reproduced using provided checkpoints. denotes inhouse retraining using published codes. Grayed rows in (c) represent methods that finetune the encoder. Top performances for each metric and dataset are in bold.
@@
RepNet [16] R2D50 - 0.995 - 0.013
TransRAC [20] VSwinT  9.130 0.443  0.085 0.291
MFL [30$\dagger$ VSwinT - 0.384 - 0.386
ESCounts VSwinT 6.905 0.298  0.1830.183\;0.183\;0.183 0.403
ESCounts VMAE  $\;\mathbf{4.455}\;$ $\mathbf{0.213}$  $\;\mathbf{0.245}\;$ $\mathbf{0.563}$
@@
Sight & Sound [67$\dagger$ R(2+1)D18 - 0.307 - 0.511
@@
ESCounts VMAE $\mathbf{3.029}$ $\mathbf{0.276}$ $\mathbf{0.319}$ 0.673
@@
TransRAC [20] VSwinT - 0.640 - 0.324
MFL[30$\dagger$ RX3D101 - 0.388 - 0.510
@@
Sight & Sound [67$\dagger$ R(2+1)D18 - $\mathbf{0.143}$ $\mathbf{0.800}$
Table 2: Cross-dataset generalisation scores. Arrows $X\rightarrow Y$ denote train dataset $X$ and test dataset $Y$. Results obtained using provided checkpoints are denoted with .
RepCount $\rightarrow$ UCFRep RepCount $\rightarrow$ Countix
@@
RN [16] - 0.998 - 0.009 - - - -
TRAC [20] 6.701 0.640 0.087 0.324 6.867 0.593 0.132 0.364
MFL [30] - 0.523 - 0.350 - - - -
ESCounts $\mathbf{3.536}$ $\mathbf{0.317}$ $\mathbf{0.219}$ $\mathbf{0.571}$ $\mathbf{4.429}$ $\mathbf{0.374}$ $\mathbf{0.185}$ $\mathbf{0.521}$
Tab. 2 - RepCount. We demonstrate that our ESCounts model outperforms recent methods [20, 30] with the same Swin Tiny (VSwinT) [35] encoder, by $-0.15$ in MAE and $+0.11$ in OBO. Our approach also improves upon [30] which uses optical flow and video in tandem, by margins of $-0.09$ in MAE and $+0.02$ in OBO. The largest improvements are observed with VideoMAE (VMAE) surpassing [20] on the introduced OBZ by $+0.16$ and reducing RMSE by $-4.68$.
@@
Tab. 2 - Countix. Compared to the state-of-the-art [16, 67] our ESCounts consistently outperforms other models with the same R(2+1)D18 encoder. Our video-only model surpasses the audio-visual model in [67] by $+0.19$ OBO. Further improvements on the RMSE, MAE, and OBZ are observed with VMAE.
@@
Tab. 2 - UCFRep. Across approaches with frozen encoders ESCounts improve the top-performing [30] by $+0.19$ OBO and $-0.17$ MAE. Our method does not outperform [67, 65] that fine-tune their encoders on the training set of UCFRep which is advantageous given the dataset’s size as also noted in [30]. We showcase this with Context (F) trained from the publicly available code of [65] with a frozen encoder and result in a significant drop in performance with +0.51 MAE and -0.42 OBO. Despite this, we achieve strong results.
@@
Figure 5: Scatter plot, instances, and density maps from RepCount, Countix and UCFRep. The dotted diagonal denotes correct predictions. We compare ESCounts against TransRAC [20] on Repcount and Context [65] on UCFRep. Examples a--m are marked with corresponding action class and count predictions. We add the Ground Truth (GT) and Predicted (P) density maps for each instance. For Countix, pseudo-labels are shown as GT, as no start-end times are given.
@@
Qualitative Results. In Fig. 5 we visualise predicted to ground truth counts. For RepCount and UCFRep, we select [20] and [65] as respective baselines and use their publicly available checkpoints444[30] was not used as a baseline as the code is not public. The publicly available checkpoints on Countix obtained lower results than originally reported.. ESCounts accurately predicts the number of repetitions for a wide range of counts, with most predictions being close to the diagonal. Though predictions from both the baseline and ESCounts are close to the ground truth in low counts, they significantly diverge in high counts. We additionally visualise specific examples with sample frames from these videos. ESCounts is robust to the magnitude of counts, with accurate predictions over low count (a,b,g,k,l) and high count (d,i,m) examples. In cases of over- and under-predictions; e.g. (c,e,f,h,j,n) ESCounts predictions remain closer to actual counts. As shown by the density maps, ESCounts can also localise the repetitions. For Countix, because the model was trained on pseudo labels, even though ESCounts can predict accurate counts, it struggles to localise the repetitions at times. We further investigate the localisation capabilities in the supplementary.
@@
Cross-dataset Generalisation. Following prior works [20, 30], we also test the generalisation capabilities of our method. We use our model trained on RepCount and evaluate it on the test sets of UCFRep and Countix. As shown in Tab. 2, our method achieves the best results surpassing [30] by $-0.21$ in MAE and $+0.22$ in OBO in UCFRep. We also explore cross-dataset generalisation for Countix, and select [20] as a baseline with the public checkpoint. We outperform [20] by significant margins across metrics, showcasing the strong ability of ESCounts to generalise across datasets.
@@ 4.3 Ablation Studies
In this section, we conduct ablation studies on RepCount [20] using VMAE as the encoder. We study the impact of exemplars by replacing cross- with self-attention and vary the number of exemplars used during training. We evaluate the sensitivity of our method to the exemplar sampling probability $p$. We ablate $\sigma$ in density map estimation and analyse the impact of time-shift augmentations and our loss function’s components.
@@
$|\mathcal{S}|\sim\{0,1,2\}$ $\mathbf{4.455}$ $\mathbf{0.213}$ $\mathbf{0.245}$ $\mathbf{0.563}$
@@
Do exemplars help in training? We study the impact of using exemplars for training by directly replacing the cross-attention decoder blocks with self-attention. As seen in Tab. 3, using self-attention (SA-only) performs significantly worse than our proposed ESCounts. Cross-attending exemplars decrease the RMSE/MAE by $-1.20$ and $-0.06$ whilst improving OBZ and OBO by $+0.10$ and $+0.09$, thus emphasising the benefits of exemplar-based VRC.
How many exemplars to sample? Since we use a varying number of exemplars whilst training, we also experiment by varying $|\mathcal{S}|$ from $0-2$ in Tab. 3. In $|\mathcal{S}|=0$ we train only with the learnt zero-shot latent. Training with $|\mathcal{S}|\sim\{0,1,2\}$ provides the best zero-shot scores in inference. We believe training with a larger and more diverse set of exemplars helps the model to learn a general zero-shot token that can capture general information about repetitions. Due to memory constraints, we could not train with more than 2 exemplars.
How to sample exemplars? We analyse the impact of sampling exemplars from other videos, from the same or a different action class. As expected, Tab. 3 shows that keeping the same action category for both exemplars and query video performs the best, as ensuring the same action semantics between exemplars and query video helps to learn their correspondence. In this table, we used sampling probability $p=0.4$.
We vary the sampling probability from other videos of the same underlying action $p$ in Tab. 3. For $p=0.0$ exemplars are sampled exclusively from the query video, whilst for $p=1.0$ exemplars are sampled solely from other videos of the same class. The best performance was observed with $p=0.4$, showcasing sampling from the same video to be critical for learning.
@@
What’s the impact of time-shift augmentations? As predictions are aggregated over $|\mathbf{K}|$ density maps by time-shifting the video input. As shown in Tab. 3, having $|\mathbf{K}|=4$ shifted start/end positions provides the best results. However, results are strong even without test-time augmentations in $|\mathbf{K}|=1$.
@@
How close are predictions to the ground truth?. We further relax the off-by metrics to Off-By-N in Figure 6(a) to visualise the proximity of predictions to the ground truth. Overall, 84% of predictions are within ±$\pm 3$ of the actual count.
What is the performance per action category?. In Figure 6(b), we plot the OBZ per action class. ESCounts performs fairly uniformly across all classes with the best performing categories being pommelhourse and squat.
@@
In Figs. 6(g), 6(h), 6(i) and 6(j) we report VRC metrics with results grouped by average video repetition duration. These are grouped, into equal sized bins, to XS=(0-0.96)s, S=(0.96-1.53)s, M=(1.53-2.29)s, L=(2.29-3.09)s, XL=(>3.09)s. Predicting density maps is more challenging for short repetitions. However, ESCounts can still correctly predict counts across repetition lengths as shown by Figures 6(i) and 6(j). We also group videos by duration into XS=(8.0-11.0)s, S=(11.0, 26.0)s, M=(26.0, 33.9)s, L=(33.9-45.9)s and XL=(45.9-68.0)s. From Figures 6(k), 6(l), 6(m) and 6(n), counting repetitions from longer videos is more challenging.
@@
4.4 Inference with More Shots
up to this point, only the learnt zero-shot latent was used for inference.
Prior works in object counting [10, 38] have reported results using exemplars (i.e. object crops) in inference. While this is not comparable to other works in VRC, we can assess our method’s ability to utilise exemplars during inference in Tab. 4. Video exemplars steadily improve performance as the number of exemplars increases. As our model cross-attends exemplars in parallel, we can train with $0-2$ exemplars, and use $>2$ exemplars during inference. We show comparable results when sampling exemplars from the test video vs exemplars from the training set videos showcasing the same action category. Combined with a classifier, one can envisage an approach that classifies the action then sources exemplars from the training set to assist counting during inference.

### v2

@@ 3.3 Time-Shift Augmentations
The predicted density map $\tilde{\mathbf{d}}$ results from encoding a video with $\mathcal{E}$, over non-overlapping sliding windows. However, as each window is of fixed temporal resolution, repetitions may span over multiple windows. Thus, we include time-shifting augmentations in which the start time of the encoded video is adjusted to allow for different spatiotemporal tokens. We train with augmentations of the start time whilst at inference, we use an ensemble of time-shift augmentations for a more robust estimation. We use multiple overlapping sequences as shown in Fig. 4 and combine the predicted density maps over $\mathbf{K}$ shifted start/end positions. We obtain the final predicted density map by temporally aligning and averaging the predictions; $\tilde{\mathbf{d}}_{t}=\frac{1}{|\mathbf{K}|}\underset{k\in\mathbf{K}}{\sum}\tilde{\mathbf{d}}^{k}_{t+\epsilon_{k}}$, where $\epsilon_{k}$ is the shifting for each $k\in\mathbf{K}$.
@@ 4 Experiments
We overview the used datasets, implementation details, and evaluation metrics in Sec. 4.1. We include quantitative and qualitative comparisons to state-of-the-art methods in Sec. 4.2. We ablate over different ESCounts settings in Sec. 4.3. For all results, we only report zero-shot counting during inference. In Sec. 4.4, we evaluate ESCounts’ when exemplars are available during inference.
@@ 4.1 Experimental Setup
RepCount [22] contains videos of workout activities with varying repetition durations. Annotations include counts alongside start and end times per repetition. We use the publicly available Part-A with $758$, $131$, and $152$ videos for train, val, and test respectively. Additionally, we use the provided open set split with 70% categories for train, 10% for val, and 20% for testing. We tune the hyperparameters on the val set and report our results on the test set.
@@
Countix [15] is a subset of Kinetics [9] containing videos of repetitive actions with $4,588$, $1,450$, and $2,719$ videos for train, val, and test respectively. Counts are provided without individual repetition start-end times. Countix does not have many pauses or interruptions between counts. Thus, we define pseudo-repetition annotations by dividing videos into uniform segments based on the ground truth count. The pseudo-labels are used to estimate the density maps without additional annotations, to compare directly to other methods.
@@
UCFRep [70] is a subset of UCF-101 [57] consisting of 420 train and 106 val videos from 23 categories with counts and annotations of start and end times. Following [70, 33], we report our results on the val split as no test set is available.
@@
Implementation Details. Unless specified otherwise, we use MAE-pretrained ViT-B [17] as our encoder $\mathcal{E}$ with Kinetics-400 [9] weights. We sample frames from variable-length videos every 4 frames using a sliding window of $64$ frames. At each window, our encoder’s input is of $16\times 224\times 224$ size, and the output is $8\times 14\times 14$, resulting in $1568$ spatiotemporal tokens. We use $C=512$ channels333when encoders have a different output, we add a fully connected layer to map to $C$. The input to the decoder is of variable length $M=1568\frac{\mathcal{R}}{64}$, where $\mathcal{R}$ is the total number of frames in the video at raw framerate. Exemplars are sampled uniformly with $16$ frames between the start and end of a repetition.
@@
The encoder is frozen and we only train the decoder and zero-shot latent $\mathbf{z}_{0}$. We use $L=2$ and $L^{\prime}=3$ ablating this choice in Sec. 7 of the appendix. We train for 300 epochs on a single Tesla V100 with a batch size of 1, to deal with variable-length videos, accumulating gradients over 8 batches. We use $5e^{-2}$ weight decay and a learning rate of $5e^{-5}$ with decay by $0.8$ every $60$ epochs. Per training instance, we randomly set the number of exemplars $|\mathcal{S}|\!\sim\!\{0,1,2\}$ and sample $\mathcal{S}$ exemplars. We set the chance of sampling exemplars from a different video to $p=0.4$.
@@
Only the learnt latent are used at inference to predict repetition counts. We aggregate predictions over $|\mathbf{K}|=4$ sequences.
@@
Evaluation Metrics. Following previous VRC works [22, 15, 72], we use Mean Absolute Error (MAE) and Off-By-One accuracy (OBO) as evaluation metrics, calculated as Eqs. 7 and 8 respectively. Inspired by image counting methods [38, 3], we introduce Root-Mean-Square-Error (RMSE) in Eq. 9 for VRC providing a more robust metric for diverse counts compared to MAE’s bias towards small counts. We also report the off-by-zero accuracy (OBZ) in Eq. 10 as a tighter metric than the corresponding OBO for precise counts.
@@ 4.2 Comparison with State-of-the-art
In Tab. 2, we compare ESCounts, to prior methods on the three datasets. We provide results on the same backbone as the best-performing method on each dataset, for fair and direct comparison to previous works.
Table 1: Comparison of VRC methods. $\dagger$ represents multi-modal models that use added audio or flow. denotes results reproduced using provided checkpoints. denotes inhouse re-training using published codes. Grayed rows in (c) represent methods that finetune the encoder. Top performances for each metric and dataset are in bold.
@@
RepNet [15] R2D50 - 0.995 - 0.013 - -
TransRAC [22] SwinT  9.130 0.443  0.085 0.291 0.625 0.204
MFL [33$\dagger$ SwinT - 0.384 - 0.386 - -
DeTRC [34] ViT-B - 0.262 - 0.543 - -
SkimFocus [73] SwinB - 0.249 - 0.517 - -
ESCounts SwinT 6.905 0.298 $0.183$ 0.403 - -
ESCounts ViT-B $\mathbf{4.455}$ $\mathbf{0.213}$ $\mathbf{0.245}$ $\mathbf{0.563}$ $\mathbf{0.436}$ $\mathbf{0.519}$
@@
Sight & Sound [72$\dagger$ R(2+1)D18 - 0.307 - 0.511
@@
ESCounts ViT-B $\mathbf{3.029}$ $\mathbf{0.276}$ $\mathbf{0.319}$ 0.673
@@
TransRAC [22] SwinT - 0.640 - 0.324
MFL[33$\dagger$ RX3D101 - 0.388 - 0.510
@@
Sight & Sound [72$\dagger$ R(2+1)D18 - $\mathbf{0.143}$ $\mathbf{0.800}$
Table 2: Cross-dataset generalisation. Arrows denote train $\rightarrow$ test datasets. Results with provided checkpoints are denoted with .
RepCount $\rightarrow$ Countix RepCount $\rightarrow$ UCFRep
@@
RepNet [15] - - - - - 0.998 - 0.009
TransRAC [22] 6.867 0.593 0.132 0.364 6.701 0.640 0.087 0.324
MFL [33] - - - - - 0.523 - 0.350
SkimFocus [73] - - - - - 0.502 - 0.391
DeTRC [34] - - - - - 0.543 - 0.418
ESCounts $\mathbf{4.429}$ $\mathbf{0.374}$ $\mathbf{0.185}$ $\mathbf{0.521}$ $\mathbf{3.536}$ $\mathbf{0.317}$ $\mathbf{0.219}$ $\mathbf{0.571}$
RepCountTab. 2 shows that ESCounts outperforms recent methods [22, 33, 34, 73]. Compared to the baseline [22], we improve OBZ by $+0.16$ and reduce RMSE by $-4.68$. We test on two backbones - SwinT [40] used in [22, 33] and ViT-B used in [34]. On the same SwinT backbone, our approach outperforms [33], which uses optical flow and video in tandem, by margins of $-0.09$ MAE and $+0.02$ OBO, showcasing ESCounts’ ability to learn repeating motions implicitly. With a ViT-B backbone, we outperform [34] by $-0.05$ MAE and $+0.02$ OBO.
We additionally compare ESCounts on the open set setting of RepCount-A, with non-overlapping action categories between train and test sets. ESCounts outperforms [22] significantly with $-0.19$ MAE and $+0.32$ OBO. Note that recent works do not report on this more challenging setup.
@@
Countix. Compared to the state-of-the-art [15, 72] in Tab. 2, our ESCounts consistently outperforms other models with the same R(2+1)D18 encoder. Our video-only model surpasses the audio-visual model in [72] by $+0.19$ OBO. Further improvements on the RMSE, MAE, and OBZ are observed with ViT-B.
@@
UCFRep. Compared to methods with frozen encoders in Tab. 2, ESCounts with ViT-B improves the previous SoTA by $+0.19$ OBO and $-0.17$ MAE and outperforms [33] on the same RX3D101 backbone by $+0.22$ OBO. Our method does not outperform [72, 70] that fine-tune their encoders on UCFRep. As noted in [33] this is advantageous given the dataset’s size. We show this experimentally by reporting Context (F) trained from the available code of [70] with a frozen encoder, resulting in a significant performance drop with +0.51 MAE and -0.42 OBO. In all directly comparable results, ESCounts achieves stronger results.
@@
Figure 5: RepCount, Countix, and UCFRep scatter plot, instances, and density maps. The dotted diagonal denotes correct predictions. We compare ESCounts against TransRAC on Repcount and Context on UCFRep. Action classes and count predictions are shown for each instance. We add the Ground Truth (GT) and Predicted (P) density maps per instance. Pseudo-labels are shown as GT for Countix.
@@
Qualitative Results. In Fig. 5 we visualise predicted to ground truth counts as scatter plots. For RepCount and UCFRep, we select [22] and [70] as respective baselines and use their publicly available checkpoints444[33] was not used as a baseline as the code is not public. The publicly available checkpoints on Countix obtained lower results than originally reported.. ESCounts accurately predicts the number of repetitions for a wide range of counts, with most predictions being close to the ground truth i.e. the diagonal. Though predictions from both the baseline and ESCounts are close to the ground truth in low counts, they significantly diverge in high counts. We visualise specific examples and their density maps. ESCounts is robust to the magnitude of counts, with accurate predictions over low (a,b,g,k,l) and high (d,i,m) count examples. In cases of over- and under-predictions; e.g. (c,e,f,h,j,n) ESCounts predictions remain closer to actual counts. As shown by the density maps, ESCounts can also localise the repetitions. For Countix, even though ESCounts can predict accurate counts, as the model was trained on pseudo labels, it struggles to localise some of the repetitions. We investigate the localisation capabilities in Sec. 8 of the appendix.
@@
Cross-dataset Generalisation. Following [22, 33], we test the generalisation capabilities of our method in Tab. 2. We use ESCounts trained on RepCount and evaluate on the Countix and UCFRep test sets. For Countix, we outperform the baseline [22] by significant margins across metrics. For UCFRep, our method surpasses [33] by $-0.21$ in MAE and $+0.22$ in OBO. ESCounts in this setting still outperforms [15, 22, 33, 34] trained on UCFRep in Tab. 2, showcasing the strong ability of ESCounts to generalise to unseen actions.
@@ 4.3 Ablation Studies
In this section, we conduct ablation studies on RepCount [22] using ViT-B as the encoder. We study the impact of exemplars by replacing cross- with self-attention and varying the number of training exemplars. We evaluate the sensitivity of our method to the exemplar sampling probability $p$, density map $\sigma$, the impact of time-shift augmentations, and the components of our objective.
@@
$|\mathcal{S}|\sim\{0,1,2\}$ $\mathbf{4.455}$ $\mathbf{0.213}$ 0.245 $\mathbf{0.563}$
$|\mathcal{S}|\sim\{0,1,2,3\}$ 4.497 0.215 $\mathbf{0.246}$ 0.560
$|\mathcal{S}|\sim\{0,1,2,3,4\}$ 4.482 0.215 0.240 0.559
@@
Do exemplars help in training? We study the impact of using exemplars for training by directly replacing the cross-attention decoder blocks with self-attention. As seen in Tab. 3, using self-attention (SA-only) performs significantly worse than our proposed ESCounts. Cross-attending exemplars decrease the RMSE/MAE by $-1.20$ and $-0.06$ whilst improving OBZ and OBO by $+0.10$ and $+0.09$, respectively. This emphasises the benefits of exemplar-based VRC.
How many exemplars to sample? A varying number of training exemplars $|\mathcal{S}|$ is used in Tab. 3. For $|\mathcal{S}|=0$, we train only the zero-shot latent $\mathbf{z}_{0}$ alongside the model’s parameters. Training with $|\mathcal{S}|\sim\{0,1,2\}$ provides the best zero-shot scores at inference with our method efficiently learning to generalise by attending to only a few exemplars. The inclusion of more exemplars saturates performance.
How to sample exemplars? In Tab. 3 we analyse the impact of sampling exemplars from the same or other training videos. As expected, keeping the same action category for both exemplar and query videos performs the best, as ensuring the same action semantics between exemplars and query video helps to learn correspondence. In this table, we used sampling probability $p=0.4$. In Tab. 3, we vary the sampling probability from other videos of the same underlying action $p$. For $p=0.0$, exemplars are sampled exclusively from the query video, whilst for $p=1.0$, exemplars are sampled solely from other videos of the same class. The best performance was observed with $p=0.4$, showcasing that the visual characteristics of exemplars from the same video are critical for VRC compared to class semantics.
@@
What’s the impact of time-shift augmentations? Predictions are aggregated over $|\mathbf{K}|$ density maps by time-shifting the video input. As shown in Tab. 3, having $|\mathbf{K}|=4$ shifted start/end positions provides the best results. However, results are strong even without test-time augmentations in $|\mathbf{K}|=1$.
@@
How close are predictions to the ground truth? We further relax the off-by metrics to Off-By-N in Fig. 6(a) to visualise the proximity of predictions to the ground truth. Overall, 84% of predictions are within ±$\pm 3$ of the actual count.
What is the performance per action category? In Fig. 6(b), we plot the OBZ per action class. ESCounts performs fairly uniformly across all classes with the best-performing categories being pommelhourse and squat.
@@
In Figs. 6(g), 6(h), 6(i) and 6(j) we report VRC metrics with results grouped by average video repetition duration. These are grouped, into equal sized bins, to XS=(0-0.96)s, S=(0.96-1.53)s, M=(1.53-2.29)s, L=(2.29-3.09)s, XL=(>3.09)s. Predicting density maps is more challenging for short repetitions. However, ESCounts can still correctly predict counts across repetition lengths as shown by Figs. 6(i) and 6(j). We also group videos by duration into XS=(8.0-11.0)s, S=(11.0, 26.0)s, M=(26.0, 33.9)s, L=(33.9-45.9)s and XL=(45.9-68.0)s. From Figs. 6(k), 6(l), 6(m) and 6(n), counting repetitions from longer videos is more challenging.
@@
4.4 Multi-Shot Inference
We use learnt latents for exemplar-free inference.
Prior object counting [38, 43] report results with exemplars (i.e. object crops) at inference. While this is not comparable to other VRC works, we can assess our method’s ability to utilise exemplars during inference in Tab. 4. Video exemplars steadily improve performance as the number of exemplars increases. Our model cross-attends exemplars in parallel, training with $0\!\!-\!\!2$ exemplars, and can even use $>\!2$ exemplars at inference. We show comparable results when sampling exemplars from the test video or training videos with the same action category. Combined with a classifier, a closed-set approach can be envisaged that classifies the action and then sources exemplars from the training set to assist counting during inference.